Q.No.1 Use vectors to decide whether the triangle with vertices P(1,-3,-2), Q(2,0,-4) and R(6,-2,-5) is right-angled.


1
Expert's answer
2020-11-24T08:38:44-0500

Let us find the coordinates of the vectors "\overrightarrow{QP}" and "\overrightarrow{QR}:"


"\overrightarrow{QP}=(1-2,-3-0,-2-(-4))=(-1,-3,2)"


"\overrightarrow{QR}=(6-2,-2-0,-5-(-4))=(4,-2,-1)"


Since the scalar product "\overrightarrow{QP}\cdot\overrightarrow{QR}=4(-1)-3(-2)+2(-1)=-4+6-2=0" , we conclude that "\angle PQR=90^\circ", and therefore the triangle with vertices "P(1,-3,-2), Q(2,0,-4)" and "R(6,-2,-5)" is right-angled.